OpenStudy (anonymous):

solve equation for y y-3=-2(x+1)^2

OpenStudy (bahrom7893):

y-3=-2(x+1)^2 add 3 to both sides: y = -2(x+1)^2 + 3

OpenStudy (anonymous):

so thats the answer?

OpenStudy (bahrom7893):

yea... i dont see anything else